stromj3688:

A battery contains two metals that have different tendencies to attract electrons. If one is lithium with an electron affinity of -3.05, and the other is zinc with an electron affinity of -0.76, describe how the electrons will flow. Then, describe how you could make this an even stronger battery.

LandonGoma:

For zinc: Zn⟶Zn++e− ΔH=9.585×102 kJ/mol Zn+⟶Zn2++e− ΔH=1.733×103 kJ/mol Zn⟶Zn2++2e− ΔH=2.692×103 kJ/mol For copper: Cu⟶Cu++e− ΔH=7.454×102 kJ/mol Cu+⟶Cu2++e− ΔH=1.958×103 kJ/mol Cu⟶Cu2++2e− ΔH=2.703×103 kJ/mol

stromj3688:

Thank you but I don't need the equations for it I need an explanation of how the electrons would flow and how could I make it a stronger battery